I have 5 questions
1) if x + 9 is an even integer, is x an even or odd integer? what is the next consecutive integer? what is the next consecutive even integer?
2) if x-1 is an odd integer what is the next consecutive odd integer? What are the next 2 consecutive even integers?
3) find four consecutive even integers such that the sum of the 2nd and 4th integers is one and a half times the sum of the first and third integers
4)find 3 consecutive even integers such that the sum of twice the 1st and three times the third is 14 more than 4 times the 2nd
5)three siblings have ages that are consecutive odd integers. find the ages if the sum of the age of the youngest and three times the age of the oldest is five less than five times the middle siblings age

Answer by CubeyThePenguin(3113)

1. x+9 is even and 9 is odd ---> x is odd

2. x-1 is odd, so the next consecutive odd integer is x+1
next 2 consecutive even numbers: x, x+2

3. consecutive even integers: x, (x+2), (x+4), (x+6)
2x + 3(x+4) = 14 + 4(x+2)
2x + 3x+12 = 14 + 4x+8
5x + 12 = 4x + 22
x = 10

The integers are 10, 12, 14, and 16.

5. consecutive odd integers: (x-2), x, (x+2)

(x-2) + 3(x+2) = 5x - 5
x-2 + 3x+6 = 5x - 5
4x + 4 = 5x - 5
9 = x

The youngest sibling is 7, the middle sibling is 9, and the oldest sibling is 11.